OpenStudy (anonymous):

why isn't the limit as x approaches -3 of (x^2 +6x + 9)/ (x^2 + 2x -3) equal to 1?

OpenStudy (anonymous):

1. Let's write it better . \[\huge \lim_{x\to (-3)}\;\;\; \frac{x^2+6x+9}{x^2+2x-3}\] 2. Let's see ! O_O

OpenStudy (anonymous):

factorizing: (x+3)(x+3)/(x+3)(x-1) = (x+3)/(x-1) taking limit equals 0

OpenStudy (anonymous):

As @myko said ... factorizing we get : \[\LARGE \lim_{x\to (-3)}\;\;\;\frac{(x+3)(x+3)}{(x-1)(x+3)}\] \[\LARGE \lim_{x\to (-3)}\;\;\;\frac{(x+3)}{(x-1)}=\frac{-3+3}{-3-1}=\frac{0}{-4}=0\] why do you think it should be equal to 1 ?

OpenStudy (anonymous):

see i didnt factor...i subsituted -3 into the numerator and denominator... i got -18/-18

OpenStudy (anonymous):

substituting we get: \[\LARGE \frac{(-3)^2+6(-3)+9}{(-3)^2+2(-3)-3}=\frac{9-18+9}{9-6-3}\] \[\LARGE \frac{-18}{9-9}=\frac{-18}{0}\] and we know that dividing by 0 is not allowed ! ... you made a mistake while calculating O_O

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

what if we factored 9 as (x- 3) (x-3)

OpenStudy (anonymous):

Glad to help... You're welcome ! ;)

OpenStudy (anonymous):

would that mattered

OpenStudy (anonymous):

how do you mean ? O_O

OpenStudy (anonymous):

((x- 3) (x-3))/ ((x-1) (x+3))

OpenStudy (anonymous):

i guess since you are trying to cancel out...you would be force to use x+3

OpenStudy (anonymous):

\[\LARGE (x-3)(x-3) \neq x^2+6x+9\] \[\LARGE (x-3)(x-3) =x^2-3x-3x+9\] \[\LARGE (x-3)(x-3) =x^2-6x+9\]

OpenStudy (anonymous):

ohhhh yeah lol thats right

OpenStudy (anonymous):

;)